I won't have time tomorrow or this weekend for an update, so here it is today.
I did not start increasing PAR yet, I wanted to make sure the polyp was stable before changing paramters. Over the last week the large polyp stayed open the majority of the time. It does close at the slightest movement, including cleaning the glass. The small polyp has NOT been seen open to date, actually I can no longer see any trace of orange color where it was.
The CB shrimp molted during the day today which I'm happy to see. On two ocassions I've seen the remains of half eaten bristle worms. I'll continue to toss the ones from my DT filter sock into the pico as I find them. I also feed 2-3 spectrum pellets daily. I was hoping to see some nitrates but not yet.
Parameters since last update
Nitrates undetectable
Ammonia undetectable
pH 7.74-8.00
Temp 77.4-79.5
ALK 7.5
CA 500
Salinty 52.8-53.1ms (1.026)
PAR 50 (due to 20K LED reading is probably up to 20% lower than actual)
Lights are on 8am to 8pm
ATO uses about 1gal evap a week
AWC still not setup, needed to get different pump heads.
Water Change 2-3 cups every 2-3 days
View attachment 12697
View attachment 12696